OpenStudy (brittany):

Write the equation in slope-intercept form of the line that has a slope of 2 and contains the point (1, 1).

OpenStudy (curry):

y= 2x-1

OpenStudy (anonymous):

y = 2x + b 1 = 2 + b b = -1 y = 2x - 1

OpenStudy (anonymous):

y - 1 = 2 (x - 1) y -1 = 2x -2 y = 2x -1

OpenStudy (brittany):

ya'll are fast!

OpenStudy (anonymous):

he's faster

OpenStudy (anonymous):

he's a pain in the retricesometimes to try and beat

OpenStudy (curry):

lol

OpenStudy (curry):

but these are very easy problems

OpenStudy (anonymous):

yea..it's just a matter of typing